A columbia river tugboat can go 12 mph in still water. It travels up the rivier 39 miles and back 30 miles in a total time of 6 hours. What is the speed of the current?
Answer by stanbon(75887)

A columbia river tugboat can go 12 mph in still water. It travels up the river 39 miles and back 30 miles in a total time of 6 hours. What is the speed of the current?
---------------------------
Upriver DATA:
distance = 39 miles ; rate = (12-c) mph; time = d/r = 39/(12-c) hrs.
---------------------------------
Downriver DATA:
distance = 30 miles ; rate + (12+c) mph ; time = d/r = 30/(12+c) hrs.
====================
Equation:
time + time = 6 hrs
39/(12-c) + 30/(12+c) = 6
----
39(12+c) + 30(12-c) = 6(12-c)(12+c)
------
829 + 9c = 864 - 6c^2
---------
6c^2 + 9c - 35 = 0
c = 1.7790 mph (speed of the current)
=========================================
Cheers,
Stan H.
==================